A contractor estimates the probabilities for the number of days required to complete a certain type of project as follows:



Time (days) 1 2 3 4 5


Probability 0.04 0.21 0.34 0.31 0.10



(4) If the contractor's project cost is made up of two parts: a fixed cost of $100 million plus $10 million for each day taken to complete the project, find the variance of total project costs. (3 Marks) 



1
Expert's answer
2021-02-01T08:04:49-0500

Solution:


First, we need to find the cost for each day, using formula: "f(x)=10x +100" , where "x" is the number of days. Calculating it for each day, we get the following distribution:


"\begin{Bmatrix}
 \xi \\
 p(\xi)
\end{Bmatrix} = \begin{Bmatrix}
 110 & 120 & 130 & 140 & 150 \\
 0.04 & 0.21 & 0.34 & 0.31 & 0.1
\end{Bmatrix}"


Remembering the formula for variance:


"var(\xi) = E(\xi^2) - (E(\xi))^2" ,


we also need to find distribution for "E(\xi^2)" :


"\begin{Bmatrix}
 \xi^2 \\
 p(\xi^2)
\end{Bmatrix} = \begin{Bmatrix}
 1210 & 1440 & 1690 & 1960 & 2250\\
 0.04 & 0.21 & 0.34 & 0.31 & 0.1
\end{Bmatrix}"


Next:


"E(\xi) = 0.04 * 110 + 0.21 * 120 + 0.34 * 130 + 0.31 * 140 + 0.1*150 = 132.2"

"E(\xi^2) = 0.04 * 12100 + 0.21 * 14400 + 0.34 * 16900 + 0.31 * 19600 + 0.1 * 22500 = 17580"


And we geat the answer:


"var(\xi) = E(\xi^2) - (E(\xi))^2 = 17580 - (132.2)^2 = 103.16"


Answer:


"var(\xi) =103.16"